Nanki Corporation purchased equipment on January 1, 2016, for $630,000. In 2016 and 2017, Nanki depreciated the asset on a straight-line basis with an estimated useful life of eight years and a $14,000 residual value. In 2018, due to changes in technology, Nanki revised the useful life to a total of 6 years with no residual value. What depreciation would Nanki record for the year 2018 on this equipment?
Answer #1

Depreciation charged under original life for 2016 and 2017 = (630000-14000)×2/8. =. 154000$

Book value at value at the beginning of 2018. =.

630000-154000. =. 476000$

Depreciation for 2018. =. 476000 × 1/4 = 119000

Depreciation that Manik record for the year 2018 is

$119000

Thank you

Hope you understand

Please be comment if you have any doubt in it